OpenStudy (anonymous):

!!! Basic Fractions, Decimals and percentages !!! !!! One Medal Question !!! !!! Fan and Testimonial !!!

OpenStudy (anonymous):

Hi Jigglypuff :)

OpenStudy (anonymous):

How can I help you?

OpenStudy (anonymous):

Hello Kirby!!!! In a school, 37% of the pupils come to school by bus, 9/25 come by car and the rest work. What percentage of the pupils walk to school?

OpenStudy (anonymous):

"KirbyLegs

OpenStudy (anonymous):

Do the rest walk or do they work?

OpenStudy (anonymous):

walk

OpenStudy (anonymous):

lol it was a mistake

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

27% walk to school

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

That is correct. Do you understand why?

OpenStudy (anonymous):

Percent is based off of 100, so you simply have to convert the given data into something divisible by 100 37% = 37/100 9/25 = 36/100

OpenStudy (anonymous):

and she's gone Xd